OpenStudy (yrivers36):

@simplify if possible (sqrt3)-1/(sqrt3)+1 explanation please -)

OpenStudy (anonymous):

\[\frac{\sqrt{3}-1}{\sqrt{3}+1} = \frac{(\sqrt{3}-1)(\sqrt{3}-1)}{(\sqrt{3}+1)(\sqrt{3}-1)} = \frac{3-2\sqrt{3}+1}{3-1} = 2 -\sqrt{3}\]

OpenStudy (yrivers36):

Thank you I got same answer but thought it was wrong because of the -